# MoC-System: Efficient Fault Tolerance for Sparse Mixture-of-Experts Model Training

### v2

@@ 6.2.3. Asynchronous Checkpointing
Table 3. Accuracy results (%) of the models on downstream tasks, pre-trained as shown in Figure 13. The downstream tasks includes: HellaSwag (zellers2019hellaswag, ), PIQA (bisk2020piqa, ), WinoGrande (sakaguchi2021winogrande, ), BoolQ (clark2019boolq, ), ARC-Easy (clark2018think, ), OpenBookQA (OpenBookQA2018, ), RACE (lai-etal-2017-race, ), MathQA (amini2019mathqa, ). “Ckpt” indicates the relative total checkpoint size compared to the baseline, which saves the full model states. “Deviation” shows the deviation of the minimum and maximum accuracy of our methods from the baseline.
@@
Method Ckpt HellaSwag PIQA WinoGrande BoolQ ARC-E OpenBookQA RACE MathQA Avg.
@@
Figure 12. Duration of a training iteration across three configurations, utilizing three checkpointing methods: baseline, the asynchronous process of baseline method, and the fully optimized asynchronous process in our MoC-System.
@@ 6.3. Impact on Model Accuracy
As shown in Figure 13 (a), applying PEC to save model weights (“W”), optimizer states (“O”), or both (“WO” and “WO-2L”) results in a validation loss curve comparable to the baseline, which saves the full states during the pre-training of the GPT-350M-16E model.
Given the similar training curves across different checkpointing methods, we further evaluate downstream tasks for each pre-trained model. Compared to the baseline method, which retains all states, our lossy methods (“W”, “O”, “WO” and “WO-2L”) achieve higher average accuracy, ranging from 0.62% to 1.08%, as shown in Table 3. Notably, our methods show the most significant accuracy improvement on the BoolQ task, ranging from 2.75% to 6.97%. We hypothesize that this level of improvement may result from state loss caused by our PEC, acting as a variant of dropout (srivastava2014dropout, ), which helps prevent overfitting in certain domains.
(a) GPT-350M-16E
(b) SwinV2-MoE
@@
We evaluate the effectiveness of our two-level PEC saving and recovery scheme in minimizing PLT and maintaining model accuracy. Given the faster speed of the snapshot process compared to the persist process, we configure $K_{persist}=1$ and experiment with varying $K_{snapshot}$ values, as depicted in Figure 14 (a). Compared with the baseline ($K_{snapshot}=1$, $K_{persist}=1$) setup, increasing $K_{snapshot}$ markedly reduces PLT, owing to the retrieval of partial experts from the in-memory snapshots on the non-fault node. Moreover, the two-level recovery with the ($K_{snapshot}=4$, $K_{persist}=1$) setup (“WO-2L” in Table3) achieves the highest average accuracy on downstream tasks, exceeding the baseline by 1.08%.
@@ 6.3.2. Sequential versus Load-aware Selection
We conduct experiments on the SwinV2-MoE model pre-training to evaluate the impact of different partial expert selection methods on model accuracy. As shown in Figure 13 (b), the three methods—baseline, PEC with sequential selection, and PEC with load-aware selection—exhibit minimal differences, with less than a 0.0012% variance in test accuracy after 80 training epochs. Considering that load-aware selection incurs additional control and synchronization costs while maintaining comparable accuracy, sequential selection appears to be the more practical choice for real-world applications. Additionally, these experiments confirm that our PEC method is applicable to both language and vision models.
@@ 6.3.3. Dynamic-K
We evaluate the efficacy of our proposed dynamic-K strategy in ensuring that the PLT does not exceed the pre-set threshold of 3.75% as the number of faults increases. As shown in Figure 14 (b), the value of $K_{pec}$ dynamically adjusts from 1 to 4, in response to escalating fault occurrences. With this strategy, the cumulative PLT remains at a low level, whereas a constant setting of $K_{pec}=1$ results in a linear increase.
(a) Tow-Level Recovery
(b) Dynamic-k
Figure 14. (a) shows the correlation between PLT and various combinations of $K_{snapshot}$ and $K_{persist}$, using tow-level recovery. The error bar represents the fluctuation in measured values. (b) demonstrates the efficacy of our Dynamic-K strategy in reducing PLT, with the red line tracking the dynamic adjustments of $K_{pec}$. These experiments are conducted during the pre-training of the GPT-350M-16E model in Case2.
@@ 7. Conclusion & Future Work
While existing LLM checkpointing ensures algorithm invariance, the MoC-System illustrates the feasibility of a more flexible algorithm-system co-designed approach to fault tolerance. As system efficiency becomes increasingly important in LLM development, more algorithms are being co-designed to enhance efficiency during training and inference. We believe fault tolerance can also be more closely integrated with LLM algorithms. In future work, we intend to explore features of LLMs, such as sparsity, to develop more efficient co-design strategies of LLM training and fault tolerance.

### v3

@@ 6.2.4. Scaling and Generalizing
Table 3. Accuracy results (%) of the models on downstream tasks, pre-trained as shown in Figure 29. The downstream tasks includes: HellaSwag (Zellers et al., 2019), PIQA (Bisk et al., 2020), WinoGrande (Sakaguchi et al., 2021), BoolQ (Clark et al., 2019), ARC-Easy (Clark et al., 2018), OBQA (Mihaylov et al., 2018), RACE (Lai et al., 2017), MathQA (Amini et al., 2019). “Ckpt” indicates the relative total checkpoint size compared to the baseline, which saves the full model states. “Deviation” shows the deviation of the minimum and maximum accuracy of our methods from the baseline.
@@
Method Ckpt HellaSwag PIQA WinoGrande BoolQ ARC-E OBQA RACE MathQA Avg. ($\uparrow$)
@@
6.2.5. Modeling and Analysis
Based on the scaling and generalizing simulations detailed in Section 6.2.4, we conclude that the identified factors influence the efficiency of the checkpointing system in two ways: (1) by affecting the duration of each iteration ($T_{F\&B}$), and (2) by affecting the time required for the snapshot ($T_{Snapshot}$). Specifically, sequence length affects only $T_{F\&B}$, whereas the number of GPUs, parallelism, hardware capability, and model size influence both $T_{F\&B}$ and $T_{Snapshot}$. Together, these factors determine the checkpoint saving overhead ($O_{save}$), ideally expressed as:
(10) $O_{save}=\begin{cases}T_{Snapshot}-T_{F\&B},&\text{if }T_{Snapshot}>T_{F\&B}\\ 0.&\text{if }T_{Snapshot}\leq T_{F\&B}\end{cases}$
Based on Equation 4 in Section 2.3 and assuming a constant failure rate denotes as $\lambda$, the number of faults can be
(11) $N_{fault}\approx\lambda I_{total}.$
The total overhead associated with the existing full checkpointing method, which saves all model states and is denoted as $O_{ckpt}^{Full}$, as well as our proposed MoC method, denoted as $O_{ckpt}^{MoC}$, can be expressed as follows:
(12) $O_{ckpt}^{Full}\approx O_{save}^{Full}\frac{I_{total}}{I_{ckpt}^{Full}}+\lambda I_{total}\left(O_{restart}+\frac{I_{ckpt}^{Full}}{2}\right),$
(13) $O_{ckpt}^{MoC}\approx O_{save}^{MoC}\frac{I_{total}}{I_{ckpt}^{MoC}}+\lambda I_{total}\left(O_{restart}+\frac{I_{ckpt}^{MoC}}{2}\right).$
$O_{save}^{Full}$ and $O_{save}^{MoC}$ represent the overhead associated with saving model states for two respective methods, while $I_{ckpt}^{Full}$ and $I_{ckpt}^{MoC}$ denote the checkpointing intervals configured for each method. To ensure that the overhead of the MoC method is less than that of the full checkpointing method, the following conditions must be met:
(14) $O_{ckpt}^{MoC}<O_{ckpt}^{Full},$
(15) $\frac{O_{save}^{MoC}}{I_{ckpt}^{MoC}}+\lambda\left(O_{restart}+\frac{I_{ckpt}^{MoC}}{2}\right)<\frac{O_{save}^{Full}}{I_{ckpt}^{Full}}+\lambda\left(O_{restart}+\frac{I_{ckpt}^{Full}}{2}\right),$
(16) $\frac{O_{save}^{MoC}}{I_{ckpt}^{MoC}}+\lambda\frac{I_{ckpt}^{MoC}}{2}<\frac{O_{save}^{Full}}{I_{ckpt}^{Full}}+\lambda\frac{I_{ckpt}^{Full}}{2}.$
Based on the condition outlined in Equation 16, we can identify two strategies to leverage the advantages of our proposed MoC design: (1) By maintaining the same checkpoint interval as the full checkpointing method, MoC can reduce overhead because it has a smaller saving overhead for each checkpointing. (2) By decreasing the checkpoint interval (more frequent checkpointing) to equalize the ratio of $O_{save}$ to $I_{ckpt}$ between MoC and the full method, MoC can still reduce the total overhead. This reduction is achieved by minimizing the lost time caused by faults, as the lost time is directly proportional to the checkpoint interval.
@@ 6.3. Impact on Model Accuracy
As shown in Figure LABEL:fig:accuracy_result1_a, applying PEC to save model weights (“W”), optimizer states (“O”), or both (“WO” and “WO-2L”) results in a validation loss curve comparable to the baseline, which saves the full states during the pre-training of the GPT-350M-16E model.
Given the similar training curves across different checkpointing methods, we further evaluate downstream tasks for each pre-trained model. Compared to the baseline method, which retains all states, our lossy methods (“W”, “O”, “WO” and “WO-2L”) achieve higher average accuracy, ranging from 0.62% to 1.08%, as shown in Table 3. Notably, our methods show the most significant accuracy improvement on the BoolQ task, ranging from 2.75% to 6.97%. We hypothesize that this level of improvement may result from state loss caused by our PEC, acting as a variant of dropout (Srivastava et al., 2014), which helps prevent overfitting in certain domains.
@@
We evaluate the effectiveness of our two-level PEC saving and recovery scheme in minimizing PLT and maintaining model accuracy. Given the faster speed of the snapshot process compared to the persist process, we configure $K_{persist}=1$ and experiment with varying $K_{snapshot}$ values, as depicted in Figure LABEL:fig:twolevel_dynamic_a. Compared with the baseline ($K_{snapshot}=1$, $K_{persist}=1$) setup, increasing $K_{snapshot}$ markedly reduces PLT, owing to the retrieval of partial experts from the in-memory snapshots on the non-fault node. Moreover, the two-level recovery with the ($K_{snapshot}=4$, $K_{persist}=1$) setup (“WO-2L” in Table3) achieves the highest average accuracy on downstream tasks, exceeding the baseline by 1.08%.
@@ 6.3.2. Sequential versus Load-aware Selection
We conduct experiments on the SwinV2-MoE model pre-training to evaluate the impact of different partial expert selection methods on model accuracy. As shown in Figure LABEL:fig:accuracy_result1_b, the three methods—baseline, PEC with sequential selection, and PEC with load-aware selection—exhibit minimal differences, with less than a 0.0012% variance in test accuracy after 80 training epochs. Considering that load-aware selection incurs additional control and synchronization costs while maintaining comparable accuracy, sequential selection appears to be the more practical choice for real-world applications. Additionally, these experiments confirm that our PEC method is applicable to both language and vision models.
@@ 6.3.3. Dynamic-K
We evaluate the efficacy of our proposed dynamic-K strategy in ensuring that the PLT does not exceed the pre-set threshold of 3.75% as the number of faults increases. As shown in Figure LABEL:fig:twolevel_dynamic_b, the value of $K_{pec}$ dynamically adjusts from 1 to 4, in response to escalating fault occurrences. With this strategy, the cumulative PLT remains at a low level, whereas a constant setting of $K_{pec}=1$ results in a linear increase.
Figure 31. (a) shows the correlation between PLT and various combinations of $K_{snapshot}$ and $K_{persist}$, using two-level recovery. The error bar represents the fluctuation in measured values. (b) demonstrates the efficacy of our Dynamic-K strategy in reducing PLT, with the red line tracking the dynamic adjustments of $K_{pec}$. These experiments are conducted during the pre-training of the GPT-350M-16E model in Case2.
Table 4. Accuracy results from fine-tuning the OLMoE (Muennighoff et al., 2024) model using various methods. “Base” refers to the pre-trained model without fine-tuning, “FT-w.o.E” indicates the fine-tuned model without fine-tuning all expert parameters, “FT-Full” represents the fine-tuned model with full state saving at each checkpointing, and “FT-PEC” denotes the fine-tuned model utilizing PEC that saves 1/8 of the experts at each checkpoint. The tasks includes: HellaSwag (Zellers et al., 2019), PIQA (Bisk et al., 2020), WG (Sakaguchi et al., 2021), BoolQ (Clark et al., 2019), ARC-C (Clark et al., 2018), OBQA (Mihaylov et al., 2018), RTE (Wang et al., 2019).
Method HS PIQA WG BQ ARC OBQA RTE Avg.
Base 57.99 80.52 68.59 74.46 47.27 44.80 54.51 61.16
FT-w.o.E 58.58 81.88 68.51 76.82 48.72 45.20 63.54 63.32
FT-Full 58.34 81.34 70.40 79.11 48.38 45.00 66.06 64.09
FT-PEC 58.78 81.45 70.24 79.17 48.23 45.00 65.58 64.06
6.3.4. Fault Tolerance during Fine-Tuning
In addition to the model’s pre-training phase, fine-tuning is another crucial stage that requires extended training periods and fault tolerance. To evaluate the impact of our proposed PEC during the fine-tuning phase, we conduct experiments using the Alpaca dataset (Taori et al., 2023) to fine-tune the open-source, pre-trained OLMoE model (Muennighoff et al., 2024). We set a fault interruption occurring halfway through the process. As shown in Table 4, PEC maintains accuracy comparable to the full-saving method. Additionally, we conduct experiments on fine-tuning with freezing all the expert parameters. This approach still achieves an increase in average accuracy, from 61.16% to 63.32%, with only a slight degradation of 0.77% compared to full-parameter fine-tuning. These results further substantiate that the expert parameters are less sensitive to a limited number of updates.
@@ 7. Conclusion & Future Work
While existing LLM checkpointing ensures algorithm invariance, the MoC-System illustrates the feasibility of a more flexible algorithm-system co-designed approach to fault tolerance. As system efficiency becomes increasingly important in LLM development, more algorithms are being co-designed to enhance efficiency during training and inference. We believe fault tolerance can also be more closely integrated with LLM algorithms. In future work, we intend to explore features of LLMs, such as sparsity, to develop more efficient co-design strategies for LLM training and fault tolerance.
Acknowledgements.
We thank the anonymous reviewers and our shepherd, Yiran Chen, for their valuable comments and suggestions. This work was supported in part by the National Key R&D Program of China (No. 2024YFB4505800), the National Natural Science Foundation of China (No. 62402411), the Guangdong Basic and Applied Basic Research Foundation (No. 2023A1515110353), the Guangdong Provincial Talent Program (No. 2023QN10X252), and the Guangzhou-HKUST(GZ) Joint Funding Program (No. 2024A03J0624). This research was conducted on the High-Performance Computing Platform of HKUST(GZ).